OpenStudy (anonymous):

How do you solve for Y? when the equation looks like this? x=y/y+1? The answer is y= x/x-1 but I don't know how they got there. Thanks

OpenStudy (anonymous):

\[x=\frac{ y }{ y+1 }\]

OpenStudy (anonymous):

answer is : \[Y=\frac{ x }{ 1-x }\]

OpenStudy (cruffo):

multiply both sides of the equation by the denominator (y+1): \[ \large (y+1)\left( x\right) = (y+1)\left(\frac{y}{y+1} \right)\] Then simplify: \[\large xy + x = y\] gather on the left all terms that have a y in them: \[\large xy - y = -x\] and factor out y: \[\large y(x-1) = -x\] Divide both sides by (x-1): \[\large y = \frac{-x}{x-1}\] Note that \(\dfrac{-a}{b}\) is the same as \(\dfrac{a}{-b}\), so a little rewritting... \[\large y = \frac{-x}{x-1} = \frac{x}{-(x-1)} = \frac{x}{-x+1} = \frac{x}{1-x}\]

OpenStudy (anonymous):

I feel dumb now! Thanks!

OpenStudy (precal):

cruffo is correct, @en28so you just need a little thing called "experience" and then you will be an expert at doing these types of problems Welcome to Open Study :)